# h FR-PHENO-2025-02 A general approach to quantum integration of cross sections in high-energy physics

### v1

@@ 5 Conclusion
Quantum technology holds the promise of solving complex problems that are classically intractable, or providing more efficient methods for tackling existing challenges than classical approaches allow. Meanwhile, the field of high-energy physics (HEP)—which seeks to uncover the fundamental laws of nature at the smallest scales—demands immense computational resources.
In this work, we investigated whether quantum computing could, in the future, help address key computational bottlenecks in HEP. Building on, and going far beyond, the work presented in Ref. [21], where MP first introduced the idea of using Quantum Monte Carlo Integration (QMCI) for cross-section calculations in HEP, we have made significant advancements in this direction.
We have developed a general approach for integrating non-trivial cross sections in HEP, in terms of decomposing the general integrand into products of constituent building blocks. This work leverages the Fourier Quantum Monte Carlo Integration method implemented in Quantinuum’s QMCI engine [32], developed in part by IW, along with several other key proprietary features of the engine. Specifically, we introduced two distinct approaches for preparing relativistic Breit-Wigner distributions on quantum registers—functions that appear as one of the key building blocks of generic integrands in cross-section calculations.
@@
Our findings suggest that such applications are unlikely to be practical during the noisy intermediate-scale quantum (NISQ) era. Instead, they are expected to become viable when fault-tolerant quantum devices are available. Notably, the resource estimates we provide for fault tolerance represent loose upper bounds. We anticipate that significant improvements will be achieved in the future—and thus significantly reduced resource requirements—as the field of fault-tolerant compilation matures beyond its infancy.
@@
Another avenue for improvement concerns the handling of experimental constraints, as discussed in Section 2.1, as these could be considered in further generality. Restricted integration domains based on constraints such as sums, maximum/minimum etc., can be easily implemented via the QMCI engine’s enhanced P-builder functionality. However, in more general cases, the cut functions may be much more complex, or even non-analytical, and this thus requires further exploration to ensure efficient integration for these cases. Similarly, extending the method to accommodate more general forms of cross-section integrands would broaden its applicability. Accurate preparation of relativistic Breit-Wigner distributions, which is crucial to this method, represents another area for further investigation. Alternative state-preparation methods that are tailored to the particular distribution could enhance both accuracy and resource efficiency, and also could be of use in general to the HEP community for other future quantum computing applications where resonances are required to be modelled. In particular, understanding in detail the systematic errors associated with state preparation is critical for performing precise integrations, and this issue warrants more detailed exploration. Finally, event sampling based on underlying distributions is a ubiquitous task in HEP. Exploring whether quantum computing and quantum integration could enhance efficiency or accuracy in this area would be of particular interest, and could provide significant advancements for the field. In summary, while this work lays a solid foundation for quantum integration in HEP, addressing these challenges will significantly expand its potential and applicability.
@@ Acknowledgements
The authors would like to thank Matthew DeCross and Julien Sorci for their careful reviewing of the manuscript. The authors thank Steven Herbert and Alexandre Krajenbrink for their support in this work and interesting discussions (and the latter for also reviewing the manuscript). The authors also want to thank Matthias Rosenkranz for interesting discussions. MP acknowledges support by the German Research Foundation (DFG) through the Research Training Group RTG2044.
@@ Appendix A Appendix
A.1 State preparation of relativistic Breit-Wigner distributions
@@ A.2 Analytical expressions
First, the analytical expression for the integral in Eq. 34 in Section 4.2 reads
@@
[…]
@@
Then, the analytical result for the additional integral in Eq. 37 in Section 4.3 reads
@@
[…]

### v2

@@ 5 Conclusion
Quantum technology holds the promise of solving complex problems that are classically intractable, or providing more efficient methods for tackling existing challenges than classical approaches allow. Meanwhile, the field of HEP—which seeks to uncover the fundamental laws of nature at the smallest scales—demands immense computational resources.
In this work, we investigated whether quantum computing could, in the future, help address key computational bottlenecks in HEP. Building on, and going far beyond, the work presented in Ref. [21], where M.P. first introduced the idea of using QMCI for cross-section calculations in HEP, we have made significant advancements in this direction.
We have developed a general approach for integrating non-trivial cross sections in HEP, in terms of decomposing the general integrand into products of constituent building blocks. This work leverages the Fourier QMCI method implemented in Quantinuum’s QMCI engine [32] (developed in part by I.W.) along with several other key proprietary features of the engine. Specifically, we introduced two distinct approaches for preparing relativistic BW distributions on quantum registers—functions that appear as one of the key building blocks of generic integrands in cross-section calculations.
@@
Our findings suggest that such applications are unlikely to be practical during the NISQ era. Instead, they are expected to become viable when fault-tolerant quantum devices are available. Notably, the resource estimates we provide for fault tolerance represent loose upper bounds. We anticipate that significant improvements will be achieved in the future—and thus significantly reduced resource requirements—as the field of fault-tolerant compilation matures beyond its infancy.
@@
Another avenue for improvement concerns the handling of experimental constraints, as discussed in Section˜2.1, as these could be considered in further generality. Restricted integration domains based on constraints such as sums, maximum/minimum etc., can be easily implemented via the QMCI engine’s enhanced P-builder functionality. However, in more general cases, the cut functions may be much more complex, or even non-analytical, and this thus requires further exploration to ensure efficient integration for these cases. Similarly, extending the method to accommodate more general forms of cross-section integrands would broaden its applicability. Accurate preparation of relativistic BW distributions, which is crucial to this method, represents another area for further investigation. Alternative state-preparation methods that are tailored to the particular distribution could enhance both accuracy and resource efficiency, and also could be of use in general to the HEP community for other future quantum computing applications where resonances are required to be modelled. In particular, understanding in detail the systematic errors associated with state preparation on the final estimation error is critical for performing precise integrations, and this issue warrants more detailed exploration. Finally, event sampling based on underlying distributions is a ubiquitous task in HEP. Exploring whether quantum computing and quantum integration could enhance efficiency or accuracy in this area would be of particular interest, and could provide significant advancements for the field. In summary, while this work lays a solid foundation for quantum integration in HEP, addressing these challenges will significantly expand its potential and applicability.
@@ Acknowledgements
The authors would like to thank Matthew DeCross and Julien Sorci for their careful reviewing of the manuscript. The authors thank Steven Herbert and Alexandre Krajenbrink for their support in this work and interesting discussions (and the latter for also reviewing the manuscript). The authors also want to thank Matthias Rosenkranz for interesting discussions. M.P. acknowledges support by the German Research Foundation (DFG) through the Research Training Group RTG2044.
@@ Appendix A Appendix
A.1 State preparation of relativistic BW distributions
@@ A.2 Analytical expressions
First, the analytical expression for the integral in Eq.˜34 in Section˜4.2 reads
@@
$={}$ $\frac{1}{6\Gamma_{\text{W}}M_{\text{W}}}\left[\Gamma_{\text{W}}M_{\text{W}}\left(\Gamma_{\text{W}}^{2}M_{\text{W}}^{2}-3\left(M_{\text{W}}^{2}-s\right)\left(M_{\tau}^{2}+M_{\text{W}}^{2}-s\right)\right)\log\left[\frac{\Gamma_{\text{W}}^{2}M_{\text{W}}^{2}+\left(M_{\text{W}}^{2}-s\right)^{2}}{M_{\text{W}}^{2}\left(\Gamma_{\text{W}}^{2}+M_{\text{W}}^{2}\right)}\right]\right$.
@@
Then, the analytical result for the additional integral in Eq.˜37 in Section˜4.3 reads
@@
$={}$ $\frac{M_{\tau}^{2}}{4\Gamma_{\text{W}}}\bigg{[}-2M_{\text{W}}\left(\Gamma_{\text{W}}^{2}\left(2s-3M_{\text{W}}^{2}\right)+\left(M_{\text{W}}^{2}-s\right)^{2}\right)\left(\cot^{-1}\left(\frac{\Gamma_{\text{W}}M_{\text{W}}}{M_{\text{W}}^{2}-s}\right)-\tan^{-1}\left(\frac{M_{\text{W}}}{\Gamma_{\text{W}}}\right)\right)$